# Maximum Preference Optimization with Importance Sampling

### 27 Dec 2023 v1

@@ Abstract
Preference learning is a key technology for aligning language models with human values. Reinforcement Learning from Human Feedback (RLHF) is a model based algorithm to optimize preference learning, which first fitting a reward model for preference score, and then optimizing generating policy with on-policy PPO algorithm to maximize the reward. The processing of RLHF is complex, time-consuming and unstable. Direct Preference Optimization (DPO) algorithm using off-policy algorithm to direct optimize generating policy and eliminating the need for reward model, which is data efficient and stable. DPO use Bradley-Terry model and log-loss which leads to over-fitting to the preference data at the expense of ignoring KL-regularization term when preference near deterministic. IPO uses a root-finding pairwise MSE loss to solve the ignoring KL-regularization problem, and learning an optimal policy. But IPO’s pairwise loss still can’t s make the KL-regularization to work. In this paper, we design a simple and intuitive off-policy preferences optimization algorithm from an importance sampling view, and add an off-policy KL-regularization term which makes KL-regularization truly effective. To simplify the learning process and save memory usage, we can generate regularization data in advance, which eliminate the needs for both reward model and reference policy in the stage of optimization.
@@ 1 Introduction
Reinforcement learning form human feedback (RLHF)Ouyang et al. (2022) uses reward-model-based reinforcement learning algorithm to learn the optimal policy. It first learns a reward model from the preference data, then using a on-policy PPO Schulman et al. (2017) algorithm to maximize the learned reward. The reward is learned to use Bradley-Terry model Bradley and Terry (1952), which assumes the preference score can be approximated from substituted with point-wise reward. This assumption may lead to an approximation error. The PPO algorithm is used on data sampled from generating policy, which may has a different support or distribution drift from preference pairwise data. So, RLHF also assumes reward model has an out-of-distribution generality. Additionally, the process to train reward model and on-policy PPO algorithm is complex, time-consuming, and unstable.
Direct preference optimization (DPO)Rafailov et al. (2023) combines off-policy algorithm and Bradley-Terry model to direct learn the generating policy from preference data. The off-policy algorithm based on KL-regularization reward maximization from off-RL community, which is data efficient, stable and eliminating the need for a reward model. DPO uses the implicit reward from off-policy algorithm to optimize the Bradley-Terry model. Because the Bradley-Terry model is used to approximate the preference score, so DPO is actually do preference score matching but not reward maximizing, which makes DPO suboptimal. Also, DPO uses Bradley-Terry model after KL-regularization, it leads to ignore the KL-regularization term and over-fitting the preference dataset.
Identity mapping preference optimization (IPO)Azar et al. (2023) also uses off-policy algorithm with KL-regularization to learn the generating policy from preference data. It learns to maximize preference probability under KL-regularization, which change DPO’s preference matching optimization to preference probability maximization, and use root finding mean square error(MSE) loss to solve the maximization problem(under KL-regularization). But it doesn’t know what does preference probability maximization actually learns theoretically, and because of the distribution or support shift between preference data distribution and policy generated data distribution, the KL-regularization may also fail.
In this paper, we design a simple and intuitive off-policy preferences optimization algorithm from an importance sampling view, and add an off-policy KL-regularization term which makes KL-regularization truly effective. Our key contribution of this paper can summary as follows:
@@
  • figure out KL-regularization fails using pairwise preference loss, and design an off-policy sample loss to make KL-regularation truly effective
@@
Figure 1: Maximum Preference Optimization (MPO) direct optimize preference maximization on preference data using off-policy algorithm, and use regularization data generated in advance to make KL-regularation truly effective, which also eliminate the needs for both reward model and reference policy in optimization stage.
@@ 2 Preliminaries
Preference learning typically started with a pretrained LLMs or LLMs fine-tuned on high quality data using maximum likelihood estimation. We define the final policy after this phase as $\pi_{\textnormal{ref}}$, and the data distribution $\pi_{\textnormal{ref}}$ to approximate as $x_{s},y_{s}\sim\mathcal{D}_{s}$, so
$\pi_{\textnormal{ref}}\approx\arg\max_{\pi}\mathbb{E}_{x_{s},y_{s}\sim\mathcal{D}_{s}}\log\pi(x_{s})\log(y_{s}|x_{s})$ (1)
@@
After pretraining and SFT phase, $\pi_{\textnormal{ref}}$ is prompted by context $x$, and generate two responses $y_{w},y_{l}\sim\pi_{\textnormal{ref}}(\cdot|x)$. Then $x,y_{w},y_{l}$ is labeled by humans to judge which response is preferred, and denote $y_{w}\succ y_{l}|x$ if $y_{w}$ is preferred, and $y_{l}\succ y_{w}|x$ if $y_{l}$ is preferred. We define a new symbol $I=\mathbb{I}[y_{w}\succ y_{l}|x]$, and all $<x,y_{w},y_{l},I>$ consist the preference dataset $\mathcal{D}^{p}$:
@@
At the final phase, a point-wise reward is learned explicitly or implicitly from the preference data, and a on-policy or off-policy learning algorithm is used to optimize the reward. In previous research, the point-wise reward is learned to use Bradley-Terry model. Given context x, define $r^{*}(y|x)$ as the reward of generating $y$. Bradley-Terry model assumes the preference probability $p^{*}(y_{w}\succ y_{l}|x)$ as:
where $\sigma(\cdot)$ is the sigmoid function. The $r^{*}(y|x)$ can be estimated by a parameterized model $r_{\phi}(y|x)$ via maximum likelihood estimation.
$\max_{\pi_{\theta}}\underset{{y\sim\pi_{\theta}(\cdot|x)}}{\mathbb{E}}[r_{\phi}(y|x)]-\beta\mathbb{D}_{\textnormal{KL}}[\pi_{\theta}||\pi_{\textnormal{ref}}]$ (8)
where $\mathbb{D}_{\textnormal{KL}}$ is the KL-divergence and $\beta$ is the regularization weight. This objective is optimized by on-policy REINFORCE Mnih et al. (2016) or PPO algorithm.
In this paper, we define a new pair-wise reward $r^{p}(y_{w}\succ y_{l}|x)=p^{*}(y_{w}\succ y_{l}|x)$, and design new algorithm to direct optimize the pair-wise reward maximum objective.
@@ 3.1 Reinforcement Learning from Human Feedback (RLHF)
The loss 9 is doing maximize likelihood estimation, and the estimated reward $r_{\phi}$ is used to approximate probability $p(y_{w}\succ y_{l}|x)$ from preference data distribution $\mathcal{D}^{p}$.
@@
The second phase of RLHF is optimizing objective 8 using $r_{\phi}$ learned from 9. PPO is an on-policy algorithm which continues collect data from the current policy $\pi_{\theta}$and estimated reward model, then it uses these data to estimate the gradient of 8, and then update the current policy. Because $\pi_{\theta}$ is different from $\pi_{\textnormal{ref}}$, samples generated by $\pi_{\theta}$ has a different distribution of $D_{\textnormal{ref}}$. So, RLHF assumes $r_{\phi}$ can generalize to out-of-distribution samples generated by $\pi_{\theta}$. Following prior work, it is straightforward to show that the optimal solution of 8 takes the form:
$\pi_{\theta}\propto\pi_{\textnormal{ref}}\exp(\frac{1}{\beta}r_{\phi}(x,y)).$ (10)
RLHF needs to learn a reward model, and maximize the learned reward using on-policy PPO algorithm, this process is complex, time-consuming, and unstable.
@@ 3.2 Direct Preference Optimization (DPO)
$r(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\textnormal{ref}}(y|x)}+\beta\log Z(x),$ (11)
where $Z(x)=\sum_{y}\pi_{\textnormal{ref}}(y|x)\exp(\frac{1}{\beta}r(x,y))$ is the partition function. Substituting the re-parameterized reward in 11 int to the Bradley-Terry model 7:
$\begin{split}p_{\theta}(y_{w}\succ y_{l}|x)&=\sigma(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)})\\ &=\frac{1}{1+\exp\big{(}\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}\big{)}}.\end{split}$ (12)
Substituting the probability 12 to loss 9, DPO formulates a maximum likelihood objective for the parameterized policy $\pi_{\theta}$ given the empirical preference dataset $\mathcal{D}^{p}$:
$\begin{split}\mathcal{L}_{\textnormal{DPO}}(\pi_{\theta};\pi_{\textnormal{ref}})=-\underset{x,y_{w},y_{l},I\sim\mathcal{D}^{p}}{\mathbb{E}}\big{[}I\cdot&\sigma(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)})\\ +(1-I)\cdot&\sigma(\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}-\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)})\big{]}.\end{split}$ (13)
In its population form, the loss takes on the form
$\begin{split}-\underset{x\sim\rho\atop y_{w},y_{l},I\sim\mu(\cdot|x)}{\mathbb{E}}\big{[}p^{*}(y_{w}\succ y_{l}|x)&\log\sigma(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)})\\ +(p^{*}(y_{l}\succ y_{w}|x))&\log\sigma(\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}-\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}))\big{]}.\end{split}$ (14)
Although this pair-wise loss eliminates the need to calculate the partition $Z(x)$, it also makes the optimal solution $\pi^{*}_{\theta}$ undefined. For example, if weight $\pi_{\theta}(y_{w}|x)$ and $\pi_{\theta}(y_{l}|x)$ with the same multiplier $M$, the logits of the sigmoid function $\sigma$ will remain the same:
$\beta\log\frac{\pi_{\theta}(y_{w}|x)*M}{\pi_{\textnormal{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)*M}{\pi_{\textnormal{ref}}(y_{l}|x)}=\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}.$ (15)
@@ 3.3 $\Psi$-PO with dientity mapping (IPO)
$\max\limits_{\pi_{\theta}}\underset{\begin{subarray}{c}x\sim\rho\\ y_{w}\sim\pi_{\theta}(\cdot|x)\\ y_{l}\sim\mu(\cdot|x)\end{subarray}}{\mathbb{E}}[\Psi(p^{*}(y_{w}\succ y_{l}|x))]-\tau\mathbb{D}_{\textnormal{KL}}[\pi_{\theta}||\pi_{\textnormal{ref}}],$ (16)
@@
$\max\limits_{\pi_{\theta}}p^{*}(\pi\succ\mu)-\tau\mathbb{D}_{\textnormal{KL}}[\pi_{\theta}||\pi_{\textnormal{ref}}].$ (17)
To optimize the objective, IPO derives an off-policy expression:
$\min_{\pi_{\theta}}\underset{x,y_{w},y_{l},I\sim\mathcal{D}^{p}}{\mathbb{E}}\big{(}h_{\pi_{\theta}}(x,y_{w},y_{l},I)-\frac{\tau^{-1}}{2}\big{)}^{2},$ (18)
where $h_{\pi_{\theta}}$ is defined as:
$h_{\pi_{\theta}}(x,y_{w},y_{l},I)=I\cdot\log\big{(}\frac{\pi_{\theta}(y_{w}|x)\pi_{\textnormal{ref}}(y_{l}|x)}{\pi_{\theta}(y_{l}|x)\pi_{\textnormal{ref}}(y_{w}|x)}\big{)}+(1-I)\cdot\log\big{(}\frac{\pi_{\theta}(y_{l}|x)\pi_{\textnormal{ref}}(y_{w}|x)}{\pi_{\theta}(y_{w}|x)\pi_{\textnormal{ref}}(y_{l}|x)}\big{)}.$ (19)
IPO claims when preferences are deterministic or near deterministic, DPO will lead over-fitting to the preference dataset at the expense of ignoring the KL-regularation term. And IPO’s loss will always regularizes $\pi_{\theta}$ towards $\pi_{\textnormal{ref}}$ by controlling the gap between the log-likelihood ratios $\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}$ and $\log\frac{\pi_{\textnormal{ref}}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}$.
Similar to DPO, the IPO loss controls the ratio of $\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}$. So the optimal policy may be undefined, and also fails the KL-regularation term.

### 31 Dec 2023 v2

@@ Abstract
Preference learning is a key technology for aligning language models with human values. Reinforcement Learning from Human Feedback (RLHF) is a model based algorithm to optimize preference learning, which first fitting a reward model for preference score, and then optimizing generating policy with on-policy PPO algorithm to maximize the reward. The processing of RLHF is complex, time-consuming and unstable. Direct Preference Optimization (DPO) algorithm using off-policy algorithm to direct optimize generating policy and eliminating the need for reward model, which is data efficient and stable. DPO use Bradley-Terry model and log-loss which leads to over-fitting to the preference data at the expense of ignoring KL-regularization term when preference is deterministic. IPO uses a root-finding MSE loss to solve the ignoring KL-regularization problem. In this paper, we’ll figure out, although IPO fix the problem when preference is deterministic, but both DPO and IPO fails the KL-regularization term because the support of preference distribution not equal to reference distribution. Then, we design a simple and intuitive off-policy preference optimization algorithm from an importance sampling view, which we call Maximum Preference Optimization (MPO), and add off-policy KL-regularization terms which makes KL-regularization truly effective. The objective of MPO bears resemblance to RLHF’s objective, and likes IPO, MPO is off-policy. So, MPO attains the best of both worlds. To simplify the learning process and save memory usage, MPO eliminates the needs for both reward model and reference policy.
@@ 1 Introduction
Reinforcement learning form human feedback (RLHF)Ouyang et al. (2022) uses reward-model-based reinforcement learning algorithm to learn the optimal policy. It first learns a reward model from the preference data, then using an on-policy PPO Schulman et al. (2017) algorithm to maximize the learned reward. The reward is learned to use Bradley-Terry model Bradley and Terry (1952), which assumes the preference score can be approximated from substituted with point-wise reward. This assumption may lead to an approximation error when preference is deterministic. The PPO algorithm is used on data sampled from generating policy, which may has a different support or distribution drift from preference data, the learned reward model inference on the out-of-distribution data may reduce the accuracy. The process of RLHF need to train reward model and on-policy PPO algorithm which is complex, time-consuming, and unstable.
Direct preference optimization (DPO)Rafailov et al. (2023) combines off-policy algorithm and Bradley-Terry model to direct learn the generating policy from preference data. The off-policy algorithm based on KL-regularization reward maximization from off-RL community, which is data efficient, stable and eliminating the need for a reward model. When preference is deterministic which occurs in most cases, the reward of Bradley-Terry model is undefined, it leads to ignore the KL-regularization term and over-fitting the preference dataset.
Identity mapping preference optimization (IPO)Azar et al. (2023) also uses off-policy algorithm with KL-regularization to learn the generating policy from preference data. It learns to maximize preference probability under KL-regularization, and use root finding mean square error(MSE) loss to solve the maximization problem and fix the KL-regularization ignorance problem. But because of the support of preference data distribution is different from reference policy distribution, both DPO’s and IPO’s KL-regularization term fails.
In this paper, we design a simple and intuitive off-policy maximum preference optimization (MPO) algorithm from an importance sampling view, and add an off-policy KL-regularization term which makes KL-regularization truly effective. Our key contribution of this paper can summary as follows:
@@
  • figure out KL-regularization fails when optimized on preference data, and design an off-policy sample loss to make KL-regularation truly effective
  • eliminate the reward substitution assumption and out-of-distribution generalization assumption
@@
Figure 1: Maximum Preference Optimization (MPO) direct optimize preference maximization on preference data using off-policy algorithm, and use offline SFT, pretrain data to make KL-regularation truly effective, which also eliminate the needs for both reward model and reference policy.
@@ 2 Preliminaries
Preference learning typically started with a pretrained LLMs or LLMs fine-tuned on high quality data using maximum likelihood estimation. We define the final policy after this phase as $\pi_{\text{ref}}$, and the data to train $\pi_{\text{ref}}$ as $\mathcal{D}_{\text{ref}}$, so,
$\pi_{\text{ref}}\approx\arg\max_{\pi}\mathbb{E}_{x,y\sim\mathcal{D}_{\text{ref}}}\log\pi(x)\log(y|x)$ (1)
@@
After pretraining and SFT phase, $\pi_{\text{ref}}$ is prompted by context $x$, and generate two responses $y_{w},y_{l}\sim\pi_{\textnormal{ref}}(\cdot|x)$. Then $x,y_{w},y_{l}$ is labeled by humans to judge which response is preferred, and denote $y_{w}\succ y_{l}|x$ if $y_{w}$ is preferred, and $y_{l}\succ y_{w}|x$ if $y_{l}$ is preferred. We define a new symbol $I=\mathbb{I}[y_{w}\succ y_{l}|x]$, and all $<x,y_{w},y_{l},I>$ consist the preference dataset $\mathcal{D}^{p}$:
@@
At the final phase, the prevailing method is using reinforcement learning algorithm to learn an explicit or implicit reward from the preference data, and then using on-policy or off-policy policy gradient algorithm to maximize the reward. Recently, some methods derive the optimal policy using reward maximization under KL-regularization and also derive a loss with optimal policy as its solution, then learn the optimal policy by minimizing the derived loss on empirical dataset.
@@ 3.1 Reinforcement Learning from Human Feedback (RLHF)
In this paper, we define a new pair-wise preference reward $r^{p}(y_{w}\succ y_{l}|x)=p^{*}(y_{w}\succ y_{l}|x)$, and design new algorithm to direct optimize the preference(reward) maximum objective.
@@
In previous research, the point-wise reward is learned to use Bradley-Terry model. Given context x, define $r^{*}(y|x)$ as the reward of generating $y$. Bradley-Terry model assumes the preference probability $p^{*}(y_{w}\succ y_{l}|x)$ as:
where $\sigma(\cdot)$ is the sigmoid function.
@@
The loss 8 is doing maximize likelihood estimation, and the estimated reward $r_{\phi}$ is used to approximate probability $p(y_{w}\succ y_{l}|x)$ from preference data distribution $\mathcal{D}^{p}$.
@@
$\max_{\pi_{\theta}}\underset{{y\sim\pi_{\theta}(\cdot|x)}}{\mathbb{E}}[r_{\phi}(y|x)]-\beta\mathbb{D}_{\text{KL}}[\pi_{\theta}||\pi_{\text{ref}}]$ (9)
where $\mathbb{D}_{\text{KL}}$ is the KL-divergence and $\beta$ is the regularization weight. This objective is optimized by on-policy REINFORCE Mnih et al. (2016) or PPO algorithm.
The second phase of RLHF is optimizing objective 9 using $r_{\phi}$ learned from 8. PPO is an on-policy algorithm which continues collect data from the current policy $\pi_{\theta}$ and estimated reward model, then it uses these data to estimate the gradient of 9, and then update the current policy. Because $\pi_{\theta}$ is different from $\pi_{\text{ref}}$ defined as 1, samples generated by $\pi_{\theta}$ has a different distribution of $D^{p}$. So, RLHF assumes $r_{\phi}$ can generalize to out-of-distribution samples generated by $\pi_{\theta}$. Following prior work Rafailov et al. (2023), it is straightforward to show that the optimal solution of 9 takes the form:
$\pi_{\theta}\propto\pi_{\text{ref}}\exp(\frac{1}{\beta}r_{\phi}(x,y)).$ (10)
RLHF also mixing a pretraining gradient into the PPO objective, in order to fix the performance regression on public NLP datasets. And RLHF call the final objective ’PPO-ptx’. Define $\mathcal{D}_{\text{pretrain}}$ as the pretraining dataset, then the combined objective defined as:
$\max_{\pi_{\theta}}\underset{\langle x,y\rangle\sim\pi_{\theta}}{\mathbb{E}}\bigl{[}r_{\theta}(x,y)]-\beta\log(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)})\bigr{]}+\gamma\underset{\mathcal{D}_{\text{pretrain}}}{\mathbb{E}}[\pi_{\theta}(x)].$ (11)
@@ 3.2 Direct Preference Optimization (DPO)
$r(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}+\beta\log Z(x),$ (12)
where $Z(x)=\sum_{y}\pi_{\text{ref}}(y|x)\exp(\frac{1}{\beta}r(x,y))$ is the partition function. Substituting the re-parameterized reward in 12 int to the Bradley-Terry model 7:
$\begin{split}p_{\theta}(y_{w}\succ y_{l}|x)&=\sigma(\beta h_{\pi_{\theta}}(x,y_{w},y_{l}))\\ &=\frac{1}{1+\exp\big{(}\beta h_{\pi_{\theta}}(x,y_{w},y_{l})\big{)}}\end{split},$ (13)
where $h_{\pi_{\theta}}(x,y_{w},y_{l})$ is defined as:
$\log\frac{\pi_{\theta}(y_{w}|x)\pi_{\text{ref}}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)\pi_{\theta}(y_{l}|x)}.$ (14)
Substituting the probability 13 to the log loss 8, DPO formulates a maximum likelihood objective for the parameterized policy $\pi_{\theta}$ given the empirical preference dataset $\mathcal{D}^{p}$:
$\begin{split}\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\underset{x,y_{w},y_{l},I\sim\mathcal{D}^{p}}{\mathbb{E}}\bigl{[}I&\sigma(\beta h_{\pi_{\theta}}(x,y_{w},y_{l}))+(1-I)\sigma(\beta h_{\pi_{\theta}}(x,y_{w},y_{l}))\bigr{]}.\end{split}$ (15)
Although this pair-wise loss eliminates the need to calculate the partition $Z(x)$, it also makes the optimal solution $\pi^{*}_{\theta}$ undefined when there are not enough constraints. For example, if weight $\pi_{\theta}(y_{w}|x)$ and $\pi_{\theta}(y_{l}|x)$ with the same multiplier $M$, the logits of the sigmoid function $\sigma$ will remain the same:
$\beta\log\frac{\pi_{\theta}(y_{w}|x)*M}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)*M}{\pi_{\text{ref}}(y_{l}|x)}=\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}.$ (16)
@@ 3.3 $\Psi$-PO with dientity mapping (IPO)
$\max\limits_{\pi_{\theta}}\underset{\begin{subarray}{c}x\sim\rho\\ y_{w}\sim\pi_{\theta}(\cdot|x)\\ y_{l}\sim\mu(\cdot|x)\end{subarray}}{\mathbb{E}}[\Psi(p^{*}(y_{w}\succ y_{l}|x))]-\tau\mathbb{D}_{\text{KL}}[\pi_{\theta}||\pi_{\text{ref}}],$ (17)
@@
$\max\limits_{\pi_{\theta}}p^{*}(\pi\succ\mu)-\tau\mathbb{D}_{\text{KL}}[\pi_{\theta}||\pi_{\text{ref}}].$ (18)
To optimize the objective, IPO derives an off-policy loss on empirical dataset:
$\min_{\pi_{\theta}}\underset{x,y_{w},y_{l},I\sim\mathcal{D}^{p}}{\mathbb{E}}\bigl{[}Ih_{\pi_{\theta}}(x,y_{w},y_{l})-(1-I)h_{\pi_{\theta}}(x,y_{l},y_{w})-\frac{\tau^{-1}}{2}\bigr{]}^{2},$ (19)
IPO claims when preferences are deterministic or near deterministic, DPO will lead over-fitting to the preference dataset at the expense of ignoring the KL-regularation term. And IPO’s loss will always regularizes $\pi_{\theta}$ towards $\pi_{\text{ref}}$ by controlling the gap between the log-likelihood ratios $\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}$ and $\log\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{l}|x)}$.
Similar to DPO, the IPO loss controls the ratio of $\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}$, not $\pi_{\theta}(y_{w}|x)$ nor $\pi_{\theta}(y_{l}|x)$. When there are not enough constraints, which in all most all cases, the optimal policy is undefined, so also fails the KL-regularation term.